A rectangle is 4 times as long as it is wide. A second rectangle is 5 centimeters longer and 2 centimeters wider than the first. The area of the second rectangle is 270 square centimeters greater than the first. What are the demensions of the original rectangle?
Answer by nerdybill(7384)

A rectangle is 4 times as long as it is wide. A second rectangle is 5 centimeters longer and 2 centimeters wider than the first. The area of the second rectangle is 270 square centimeters greater than the first. What are the demensions of the original rectangle?
.
Let w = width of original rectangle
then
4w = length of original rectangle
.
(w+2)(4w+5) = w(4w) + 270 (if you don't see this, write back)
4w^2+5w+8w+10 = 4w^2 + 270
4w^2+13w+10 = 4w^2 + 270
13w+10 = 270
13w = 260
w = 20 cm (width of original rectangle)
.
Length:
4w = 4(20) = 80 cm (length of original rectangle)